OpenStudy (anonymous):

What proportion of all Americans believe the president is doing a good job? A recent poll found that 39% of 1000 randomly selected Americans approved of the way George W. Bush is handling his job as president. a) Does this mean that only 39% of all Americans approve of the way George W. Bush is handling his job as president? b) For a sample proportion, as long as the population is greater than ten times the sample size, we can approximate the standard deviation to be the square root of p times 1 minus p divided by n What is the standard deviation for this sample?

OpenStudy (anonymous):

c) What is the critical value for a 90% confidence interval? d) What is the critical value for a 95% confidence interval? e) What is the 90% confidence interval? Be sure to include the formula with your numbers plugged in. Use the model: p plus or minus z times the square root of p times 1 minus p divided by n f) What is the 95% confidence interval?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (comm.dan):

A @schmidtdancer

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (comm.dan):

What do you think it is? @schmidtdancer

OpenStudy (anonymous):

im not sure.

OpenStudy (anonymous):

@jim_thompson5910 ? how do i start with a

OpenStudy (comm.dan):

Ok

jimthompson5910 (jim_thompson5910):

you have a SRS of 1000 people There are about 300 million people in this country. Even if you exclude young children who can't vote, that's still a lot of people.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got no, this 39 percent refers to the 1000 Americans sampled.

jimthompson5910 (jim_thompson5910):

so do you think a SRS of 1000 people could represent a population of 300 million accurately?

OpenStudy (anonymous):

what about b? i got:(b) sqrt np(1-p) = sqrt(1000(0.39)0.61) = 237.9 <----

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

use sqrt(p(1-p)/n)

OpenStudy (anonymous):

so is that b?

jimthompson5910 (jim_thompson5910):

that's part of b, yes

OpenStudy (anonymous):

is b the answer i had..?

OpenStudy (anonymous):

237.9

OpenStudy (anonymous):

is c: (0.3646 , 0.4154)

jimthompson5910 (jim_thompson5910):

s = sqrt(p(1-p)/n) s = sqrt(0.39(1-0.39)/1000) s = ???

OpenStudy (anonymous):

s=0.01542400726

jimthompson5910 (jim_thompson5910):

better

OpenStudy (anonymous):

so b is: s = sqrt(0.39(1-0.39)/1000) s= 0.01542400726

OpenStudy (anonymous):

whats c? i got (0.3646 , 0.4154)

OpenStudy (anonymous):

is b and c right^^

jimthompson5910 (jim_thompson5910):

yes b is s = sqrt(0.39(1-0.39)/1000) s= 0.01542400726

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

c?

jimthompson5910 (jim_thompson5910):

c) and d) are found using that wolfram alpha calculator we used before

OpenStudy (anonymous):

ok is what i have right tho

jimthompson5910 (jim_thompson5910):

remember how we got 1.645 and 1.96 ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

well those are the critical values for the 90% and the 95% confidence intervals

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do i find c with it

jimthompson5910 (jim_thompson5910):

just something you'll have to look up in a table or use a calc (or memorize)

jimthompson5910 (jim_thompson5910):

those are the answers to C and D because that's all those parts want

OpenStudy (anonymous):

Oh ok. so c is 1.645

OpenStudy (anonymous):

d is 1.96?

OpenStudy (anonymous):

(e) (0.3646,0.4154) (f) (0.3598 , 0.4202) are these right

jimthompson5910 (jim_thompson5910):

yep you got c and d correct

jimthompson5910 (jim_thompson5910):

hold on while i check e and f

OpenStudy (anonymous):

kk

jimthompson5910 (jim_thompson5910):

you nailed both e and f, congrats

OpenStudy (anonymous):

yay! thanks!

jimthompson5910 (jim_thompson5910):

yw